# Time to Retrain? Detecting Concept Drifts in Machine Learning Systems

### v1

@@ Results.
- The performance of CDSeer. We list the results of CDSeer side-by-side to PHT and ECHO in Table II. Our evaluation shows that CDSeer improves the accuracy of the online models for all five datasets except for one, i.e., SEA. CDSeer has comparable performance with the supervised PHT and even a slightly higher model accuracy for three of the six datasets. Compared to the semi-supervised ECHO, CDSeer achieves higher model accuracy on all six datasets. In short, our CDSeer leads to comparable performance improvement on model accuracy with the supervised PHT and outperforms the SOTA semi-supervised ECHO.
In terms of the performance in detecting concept drifts, CDSeer outperforms the SOTA semi-supervised ECHO on both precision and recall and outperforms the supervised PHT in recall but has worse precision than PHT. For the three datasets with ground-truth labels, CDSeer has acceptable precision (more than 58.9%) and high recall (more than 90%). This gives us a fair estimation of CDSeer’s precision. For the three datasets with PHT-no-retraining labels, CDSeer has much higher precision and lower recall than ECHO. This is due to ECHO generating significantly more (false) alarms. CDSeer has lower precision than PHT, which is expected as PHT has access to 100% of the manual labels.
Concerning the amount of manual labeling needed (‘Lbl’ in Table II), the proposed CDSeer has the lowest label requirements, i.e., 0.6% to 0.8%. CDSeer requires the lowest percentage of labels compared to PHT and the SOTA semi-supervised ECHO, except for the dataset Sine. Our investigation shows that the 0.03% label requirement on Sine by ECHO is because Sine does not have many data points with low estimated confidence. Thus, this resulted in a few points meeting the criteria to be selected by ECHO. However, Sine is an exception, and for the remaining datasets, ECHO requires at least 48.1% of the data for manual labeling, which is too much to be practical. Furthermore, CDSeer requests a stable range of data for manual labels compared to the large range of data required by ECHO. The count of the labeling requests (which may request for one or more labels each) shows an even larger gap. For example, the ROUTER dataset represents one month’s worth of data, where 800 data points are logged each hour. During this period, ECHO makes 22,459 requests for data labels, while CDSeer makes 80 requests. Considering the number of labels, the number of requests, and the overhead associated with processing each request, this makes a significant difference to the experts’ labeling efforts.
Overall, CDSeer shows detection performance competitive to the state-of-the-art while utilizing fewer labeled data points. Depending on the data set, our method significantly reduces the number of labels used while maintaining comparable detection results and accuracy.
@@
- Sensitivity of CDSeer’s performance. Table III shows the performance of CDSeer with different configurations including sample size and window size. First, we look at how sample size impacts our detection results. Our results show that the best sample size varies for different datasets. For ROUTER, the sample size of 1,000 works best for the ROUTER dataset with 500 being a close second. Differently, the sample size of 500 is best for Sine and SEA. For memory size, we observe that a larger memory size may sacrifice recall for precision. For different datasets though, the best-performance memory size still varies. We conclude that memory size should be adjusted according to the size of the concept drift we want CDSeer to capture (i.e., higher recall). For concept drifts that occur only briefly, memory size should be smaller and vice versa. However, this might lead to decreased precision. Last, our evaluation shows the percentage of manual labeling required by CDSeer is stable for different window and sample size combinations. Note that with the default configuration, CDSeer still outperforms the SOTA. Our experiments confirm that CDSeer can be tuned to have a better performance.
Discussions on adopting CDSeer in industry settings. We discuss the adoption of CDSeer with our industry partners and conclude that the precision and recall of CDSeer are sufficient for industry deployment. In practice, it is preferred to have higher recall to be more alerted about potential concept drifts as they can be quickly verified to rule out false alarms. In addition, CDSeer’s extremely low manual labeling effort, especially in terms of the number of requests is desirable for planning and reducing efforts. From a practical perspective, the extra effort can be diverted to examine reported concept drift. Couple that with addressing experts’ preference of labeling multiple data points at their convenience and the flexibility in the points to be labeled makes CDSeer suitable for adoption.
@@
RQ3-Generality: Can CDSeer detect concept drift on pre-trained models?
@@
Method. We include two different Keras sequential neural networks for image classifications. The first model consists of three layers and the second model consists of five layers as described in Section V-C. We trained the two models using the released data and source code from Tensorflow document [87, 78]. Model 1 is trained using the complete 60,000 data points from the FM dataset and Model 2 is trained using the 25,000 points from the LMR dataset.
@@ VII Threats to Validity
Threats to internal validity. CDSeer has parameters that may impact its performance. We conducted a sensitivity experiment to assess the impact of different parameters on the performance of CDSeer. Our experiment shows that CDSeer’s memory size should be proportional to the size of the concept drift for best recall. However, the default values are sufficient for most cases.
Threats to construct validity. We describe the threats below regarding propagation and the percentage of labels required. First, we use propagation to auto-label selected data points based on clustering, which may impact CDSeer’s performance. We included additional propagation methods for comparison. In particular, we observe that $LabelSpreading$ and $LabelPropagation$ models from sklearn [84] and observed similar results. Future works can investigate more propagation methods. Second, our method has inherent randomness from the sampling process. We expect to see a fluctuating number of labels required for manual labeling between runs. To address this issue, we report an average number from five runs in this paper.
@@ VIII Conclusion
In this work, we propose CDSeer, a concept drift detection framework that can detect concept drifts with better precision and recall compared to state-of-the-art while using fewer manual labels and having fewer constraints concerning the data distribution and model architecture. CDSeer addresses industry concerns in concept drift detection and allows for its integration into existing maintenance frameworks.

### v2

@@ Results.
- The performance of CDSeer. Table II shows the results of CDSeer side-by-side to PHT and ECHO. Our evaluation shows that CDSeer improves the accuracy of the online models for all five datasets except for one, i.e., SEA. CDSeer has comparable performance with the supervised PHT, with slightly higher model accuracy for three of the six datasets. Compared to the semi-supervised ECHO, CDSeer achieves higher model accuracy on all six datasets. In short, our CDSeer leads to comparable performance improvement on model accuracy with the supervised PHT and outperforms the SOTA semi-supervised ECHO.
As for concept drift detection, CDSeer outperforms ECHO on both precision and recall and outperforms PHT in recall but has worse precision. For the three datasets with GT labels, CDSeer has acceptable precision (more than 58.9%) and high recall (more than 90%). For the three datasets with PHT-no-retraining GT labels, CDSeer has much higher precision and lower recall than ECHO. This is due to ECHO generating significantly more (false) alarms. CDSeer has lower precision than PHT, which is expected as PHT has access to all labels.
Concerning the manual labeling needed (‘Lbl’ in Table II), CDSeer requires between 0.6% and 0.8% of the data points, which corresponds to the number of data clusters as described in section IV. This shows that CDSeer’s label requirement is the lowest for all datasets except Sine. Our investigation reveals that ECHO’s 0.03% label requirement on the Sine dataset stems from a shortage of data points with low estimated confidence, leading to a very limited number of points that meet the selection criteria. However, Sine is an exception; for the remaining datasets, ECHO requires at least 48.1% of the data for manual labeling, which is impractical. Furthermore, CDSeer requests a stable range of data for manual labeling, in contrast to the large range of data required by ECHO. The count of the labeling requests (which may request for one or more labels each) shows an even larger gap. For example, the ROUTER dataset represents one month’s worth of data, where 800 data points are logged each hour. During this period, ECHO makes 22,459 requests for data labels, while CDSeer makes 80 requests. Considering the number of labels, the number of requests, and the overhead associated with processing each request, this makes a significant difference to the experts’ labeling efforts. Overall, CDSeer shows detection performance competitive to the state-of-the-art while utilizing fewer labeled data points. Depending on the dataset, our method significantly reduces the number of labels used while maintaining comparable detection results and accuracy.
@@
- Sensitivity of CDSeer’s performance. Table III shows the performance of CDSeer with different configurations. First, we look at how window size impacts our detection results. Our results show that the best window size varies for different datasets. The window size of 1,000 works best for the ROUTER dataset while the window size of 500 is best for Sine and SEA. Similar to the window size, the best-performance memory size varies for each dataset. We observe that a larger memory size may sacrifice recall for precision. Hence, memory size should be adjusted according to the size of the concept drift we want CDSeer to capture. For concept drifts that occur only briefly, memory size should be smaller and vice versa. However, this might lead to decreased precision. Lastly, our evaluation shows that the percentage of manual labeling required by CDSeer is stable for different combinations of window and memory sizes. Labeling efforts can be further reduced at the cost of lower precision, presenting opportunities to label fewer points than requested if desired. Note that with the default configuration, CDSeer still outperforms the SOTA. Our experiments confirm that CDSeer can be tuned to have an even better performance for each dataset.
Discussions on adopting CDSeer in industry settings. We discuss the adoption of CDSeer with our industry partners and conclude that the precision and recall of CDSeer are sufficient for deployment in the industry. In practice, higher recall is preferred to be more alert to potential concept drifts, as they can be quickly verified to rule out false alarms. In addition, CDSeer’s extremely low manual labeling effort, especially in terms of the number of requests, is desirable for planning and reducing labeling efforts. From a practical standpoint, the saved effort can be redirected to examine reported concept drift. Couple that with addressing experts’ preference of labeling multiple data points at their convenience and the flexibility in the points to be labeled makes CDSeer suitable for adoption.
@@
RQ3-Generality: Can CDSeer detect concept drift in pre-trained models?
@@
Method. We include two different Keras sequential neural networks for image classifications. The first model consists of three layers and the second model consists of five layers as described in Section V-B. We trained the two models using the released data and source code from Tensorflow document [87, 78]. Model 1 is trained using the complete 60,000 data points from the FM dataset and Model 2 is trained using the 25,000 points from the LMR dataset.
@@ VII Threats to Validity
Threats to internal validity. CDSeer has parameters that may impact its performance. Our sensitivity experiment to assess the impact of different parameters on CDSeer’s performance shows that CDSeer’s memory size should be proportional to the size of the concept drift for best recall. However, the default values are sufficient for most cases.
Threats to construct validity. We describe the threats below regarding propagation and the percentage of labels required. First, the learning method to propagate the labeled data points to unlabeled data points can change CDSeer’s performance. While we observed similar results from $LabelSpreading$ and $LabelPropagation$ models from sklearn [84], future works can investigate more propagation methods. Second, our method has inherent randomness from the sampling process. We expect a fluctuating number of labels required for manual labeling between runs. To address this issue, we report an average number from five runs in this paper.
@@ VIII Conclusion
We propose CDSeer, a concept drift detection framework that outperforms the state-of-the-art in precision and recall while requiring fewer manual labels and offering more flexibility in data distribution and model architecture. CDSeer addresses industry challenges and integrates easily into existing maintenance frameworks. The implementation of CDSeer and the ROUTER dataset are close-sourced. However, we provide sufficient description to re-implement the proposed method. As for evaluation, seven of the described datasets are publicly available for reproduction.